OpenStudy (anonymous):

what is a polynomial function in standard form with zeros 1,2,-2, and -3

OpenStudy (anonymous):

ax^4+bx^3+cx^2+dx+e is this what are you asking for

OpenStudy (anonymous):

i guess so.

OpenStudy (ahsome):

\[ax^4+bx^3+cx^2+dx+e\]

OpenStudy (anonymous):

but i dont have abcde, how do i find them

OpenStudy (anonymous):

a,b,c,and d =1 and e=0

OpenStudy (anonymous):

x^4+x^3+x^2+x

OpenStudy (perl):

general form is a*(x-r_1)(x-r_2)...(x-r_n), where r_1, r_2, ... r_n are roots

OpenStudy (perl):

we can choose a = 1 , or any number you want (except zero), 1 is simplest f(x) = 1* ( x - 1) ( x - 2) ( x - (-2) ) ( x - (-3) )

OpenStudy (perl):

now your job is to simplify that, and expand it

OpenStudy (anonymous):

these are my choices. x^4 + 2x^3 +7x^2 -8x + 12 x^4 + 2x^3 -7x^2 -8x + 12 x^4 - 2x^3 +7x^2 +8x + 12 x^4 + 2x^3 +7x^2 +8x + 12

OpenStudy (perl):

of course you could copy and paste it into wolfram

OpenStudy (perl):

im not allowed to give answer :/

OpenStudy (anonymous):

i tried wolfram. it gave me decimals like 1.0006783

OpenStudy (anonymous):

i understand you cant give an answer but i dont understand how abcd =1 and e=0 then it gets to the others.

OpenStudy (perl):

where are you getting abcd = 1

OpenStudy (perl):

may I see what you put into maple, how you got 1.0006783

OpenStudy (anonymous):

the polynomial is given by (x-1)(x-2)(x+2)(x+3)={(x-1)(x+3)}{(x-2)(x+2)} =(x^2+2x-3)(x^2-4). Rest you may compute?.@minimoo

OpenStudy (anonymous):

@minimoo